# Random Lévy Looptrees and Lévy Maps

### v2

@@ 5.2.   Convergence of maps
Suppose that the previous assumptions are satisfied. Then from every sequence tending to infinity, one can extract a subsequence along which the convergence in distribution
@@
holds for the uniform topology on $[0,1]^{2}$, jointly with Theorem 5.2. The limit $D$ is a random continuous pseudo-distance which satisfies:
$D(U,\cdot)\enskip\mathop{=}^{(d)}\enskip Z^{1/3}-\min Z^{1/3},$
where $U$ is sampled uniformly at random on $[0,1]$ and independently of the rest. Furthermore, along this subsequence, we have
@@
in the Gromov–Hausdorff–Prokhorov topology. Finally, if the excursion of $X$ reduces to a constant $\beta$ times the Brownian excursion, then these subsequential limits agree with $\sqrt{\beta/3}$ times the standard Brownian sphere.

### v3

@@ 6.2.   Convergence of maps
Suppose that the previous assumptions are satisfied. Then from every sequence of integers tending to infinity, one can extract a subsequence along which the convergence in distribution
@@
holds for the uniform topology on $[0,1]^{2}$, jointly with Theorem 6.2. The limit $D$ is a random continuous pseudo-distance which satisfies the identity in law:
$(D(U,t))_{t\in[0,1]}\enskip\mathop{=}^{(d)}\enskip(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})_{t\in[0,1]},$
where $U$ is sampled uniformly at random on $[0,1]$ and independently of the rest. It also satisfies almost surely the property that for every $s,t\in[0,1]$,
$\text{if}\enskip d_{\mathscr{L}(X)}(s,t)=0,\enskip\text{then}\enskip D(s,t)=0.$
Furthermore, along any such convergent subsequence, we have
@@
in the Gromov–Hausdorff–Prokhorov topology. Finally, if the excursion of $X$ reduces to a scaled Brownian excursion or stable excursion, then the maps converge in distribution without extraction to a multiple of the Brownian sphere or a stable map respectively.